Assume that France Imports Wine and they produce it domestically as well. France is a large enough country that consumption in France affects world prices. Assume that the Demand for wine in France is given by Pd = 16 - qd and the Supply of wine in France is given by Ps = qs . Furthermore, assume that the world Supply of wine bottles is a function of the quantity in France and is given by Pw = (1/3)qF where qF represents the quantity of wine in France. If there was no import tariff on wine in France, but there was free trade with other countries, what quantity of wine would be consumed in France?
Answer #1

If there is no import tariff on wine in France, but there was free trade with other countries,

France will consume where Pd = Pw

16- q = (1/3)q

16 - q = q/3

16 = q/3 + q

16 = 4/3q

q = 16*(3/4) = 12

12 units of wine would be consumed in France.